OpenStudy (anonymous):

Use the half angle formulas to come up with an exact expression for each function value below. You DO NOT have to simplify your answer. a. cos(π/8) b. tan(π/8) c. tan(π/16)

OpenStudy (anonymous):

I've finished the first two! ** π/8 is half of π/4 ** \[a) = \sqrt{\frac{ 2+\sqrt{2} }{ 4 }}\]\[b) = \sqrt{2} - 1\]

OpenStudy (psymon):

Haha, nice. Alright, let's see about #3. Gotta use it twice D:

OpenStudy (anonymous):

Use what twice? π/8?

OpenStudy (psymon):

Well I'm just starting from pi/4 for tan is all xD Because yeah, I didn't come up with what you did for tan(pi/8)

OpenStudy (anonymous):

Should I write out what I did for B. tan(π/8) ?

OpenStudy (psymon):

If you want to. but I did come up with quite a different answer.

OpenStudy (psymon):

wait....I guess yours is correct just....simplified way more than mine I suppose xD

OpenStudy (anonymous):

Oh right! It did say you do not have to simplify it... Whoops!

OpenStudy (psymon):

I'm not even sure how you simplified it that far, haha.

OpenStudy (anonymous):

lets see...

OpenStudy (psymon):

Nvm, I see it.

OpenStudy (anonymous):

\[\tan\frac{ π }{ 8 } = \tan\frac{( \frac{ π }{ 4 }) }{ 2 } \]\[\tan(\frac{ π }{ 8 }) = \frac{ 1 - \cos \frac{ π }{ 4 } }{ \sin \frac{ π }{ 4 } }\]

OpenStudy (anonymous):

oh okay! dont mind me then :)

OpenStudy (anonymous):

I was about to write it all out... lolz

OpenStudy (psymon):

I just had to check it, lol. Hmm...you started off in adirection I wouldn't know, haha. (1-cosx)/sinx?

OpenStudy (anonymous):

It just worked! My next step was:\[\frac{ 1-\frac{ \sqrt{2} }{ 2 } }{ \sqrt{2} }\]

OpenStudy (psymon):

Looks like you missed the 2 on the bottom xD Lol, yeah, not sure what ya did o.o I just memorize the formula after all this silly time of abusing the hell out of it: \[\pm \sqrt{\frac{ 1-cosx }{ 1+cosx }}\]

OpenStudy (anonymous):

Do you mean the other √2/2 ? You're right, I did!

OpenStudy (anonymous):

And I use that formula, just not for this problem in particular...

OpenStudy (psymon):

Yeah, lol. But nah, I just memorize a lot of these. So I just jotted that down and pluggedin numbers.

OpenStudy (psymon):

Okay, well, just do what you did but with your sqrt(2) - 1, lol.

OpenStudy (anonymous):

Wait, where?

OpenStudy (psymon):

Lol, nvm. If you don't need to simplify, then I guess you could just plug in pi/8 into all your formula numbers. Unless we need to get the valueof cos and sinpi/8, too. And I see where you got your formula from, lol. I found it in my textbook. I just remember it the long way I guess xD

OpenStudy (anonymous):

All it really says is "to come up with an expression for each function".

OpenStudy (psymon):

Bleh, too many stacked square roots x_x

OpenStudy (anonymous):

I just don't know how to approach this.. lolz

OpenStudy (psymon):

You need to get the value of cos(pi/8), lol. That way you can plug it into your formula xD

OpenStudy (psymon):

The answer I got is correct but stupid xD

OpenStudy (psymon):

Look at this brutality: \[\sqrt{\frac{ 2-\sqrt{\sqrt{2}+2} }{ 2+\sqrt{\sqrt{2}+2} }}\]

OpenStudy (anonymous):

oh damn. I don't like this! UGH

OpenStudy (psymon):

Without just plugging in cos(pi/8) in, you kinda have to find the value of that and plugitin, which resultsin a ton of roots xD

OpenStudy (anonymous):

I don't even feel like attempting to get an answer like that

OpenStudy (psymon):

Is it required? xd

OpenStudy (anonymous):

Kinda? I need to be able to understand/remember/solve these types of problems when school starts

OpenStudy (psymon):

True. Well, I just did the half angle formula for cos, using pi/4 as my x and that got me the value of cospi/8. Then I just plugged that directly into the formula for half angle of tangent to finally get tan(pi/16)

OpenStudy (anonymous):

Well thats just perfect! Let me try, and thanks again :)

OpenStudy (psymon):

Mhm. Because yeah, if you use x = pi/8 in order to find tan of pi/16, then this would be the plugging in: \[\sqrt{\frac{ 1-\cos \frac{ \pi }{ 8 } }{ 1+\cos \frac{ \pi }{ 8 } }}\] Since this would be the required formula anyway, there is no practical way to get that without actually finding the value of cos(pi/8) and plugging that directly in. So using thehalf angle formula for cos, using x = pi/4, I have: \[\cos \frac{ \pi }{ 8 } = \sqrt{\frac{ 1+\frac{ \sqrt{2} }{ 2 } }{ 2 }}\] So yeah, lots of subs with square roots, but then I could use that in the same spot I have cos(pi/8) in the tan half angle formula and go from there to get more roots than a tree.

OpenStudy (anonymous):

How come you used the less simplified equation for cos(π/8)? Does mine not work in the problem?

OpenStudy (psymon):

Because I didn't realize you already had it :D

OpenStudy (anonymous):

Oh okay! But would your be easier?

OpenStudy (psymon):

But even with yours you could still do one itty bitty thing :P \[\frac{ \sqrt{2+\sqrt{2}} }{ 2 }\] And probably not xD I just did it with unsimplified versions. I would use this. Id have simplified mine to this eventually anyway, I just didnt do it immediately.

OpenStudy (anonymous):

I've gotten my final answer! \[\sqrt{\frac{2-\sqrt{2+\sqrt{2}} }{ 2+\sqrt{2+\sqrt{2}} } }\]

OpenStudy (psymon):

haha, nice xD

OpenStudy (anonymous):

THANK YOU, THANK YOU, THANK YOU!

OpenStudy (psymon):

Yep, np ^_^